# Performances in solving the Bethe-Salpeter equation with the Yambo code

### v1

@@ Abstract
In this work, we analyze the performances of two different strategies in solving the structured eigenvalue problem deriving from the Bethe-Salpeter equation (BSE) in condensed matter physics. The first strategy employs direct diagonalization, while the second is based on an iterative solver. The BSE matrix is constructed with the Yambo code, and the two strategies are implemented by interfacing Yambo with the ScaLAPACK and ELPA libraries for direct diagonalization, and with the SLEPc library for the iterative approach. We consider both the hermitian (Tamm-Dancoff approximation) and pseudo-hermitian forms, addressing dense matrices of three different sizes. A description of the implementation is also provided, with details for the pseudo-hermitian case. Timing and memory utilization are analyzed on both CPU and GPU clusters. The CPU simulations are performed on a local cluster in Rome, while the GPU simulations are performed on the Leonardo HPC cluster of CINECA. Our results demonstrate that it is now feasible to handle dense BSE matrices of the order 105.
Keywords:
Eigenvalue problems Bethe-Salpeter Equation GPU
@@ 1 Introduction
where $R$ and $A$ are dense Hermitian matrices, $R=R^{\dagger}$ and $A=A^{\dagger}$, while $C$ is a dense rectangular matrix. Here $R^{\dagger}=(R^{*})^{T}$ denotes conjugate transposition. We refer to the full BSE matrix as $H^{2p}$, and to its structure as pseudo-Hermitian [3, 4, 5, 6]. The BSE structure allows for further simplifications of the eigenvalue problem, allowing more efficient solution methods beyond the standard generalized eigenvalue approach [7]. The same structure also appears in the eigenvalue problems derived within Time-Dependent Density Functional Theory (TD-DFT) and within the Random-Phase Approximation (RPA).
@@
The BSE eigenvalue problem is coded in many condensed-matter (and quantum chemistry) packages, both open-source, such as Yambo [8, 9], BerkeleyGW [2, 10], Abinit [11, 12], Exc [13], Exciting [14, 6, 15], GPAW [16, 17], Simple [18], MolGW [19] and commercial or private, such as VASP [20, 5], TurboMole [21], and Fiesta  [22, 23]. Considering also TD-DFT and RPA, the number of scientific codes implementing the eigenvalue problem of Eq. (1) becomes very large, including applications in quantum chemistry and nuclear physics. Most of them use the standard generalized eigenvalue solvers implemented in eigensolver packages. In condensed matter, most BSE applications are done within the Tamm-Dancoff approximation (TDA) by setting $C=0$, and the BSE eigenvalue problem becomes Hermitian. However, there are physical cases where going beyond the TDA is needed, thus having to deal with the full pseudo-hermitian structure [3, 4, 5, 6].
@@
For the case of optical properties, we have $A=R^{*}$, while $C=C^{T}$. The size of both $R$ and $C$ is $N\times N$. Defining $N_{\mathbf{k}}$ as the number of k-points in the full Brillouin zone, $N_{v}$ as the number of included occupied (valence) states, and $N_{c}$ as the number of unoccupied (conduction) states, $N=N_{\mathbf{k}}\,N_{v}\,N_{c}$ is the total number of optical transitions. Numerically, the most demanding step is the building of the $R$ and $C$ matrices which grows as $\mathcal{O}(N^{2})$, with a rather large prefactor proportional to the (square of the) basis-set size, e.g. squared number of G-vectors, $\mathcal{O}(N_{G}^{2})$, for plane-wave codes. This is why, with the advent of HPC machines and complex structures, a significant effort has been devoted to improve the matrix generation algorithm, taking advantage of MPI, OpenMP and GPU porting schemes such as CUDA, Fortran or OpenACC. Thanks to these developments, matrices with a size of the order $N\propto 10^{5}$ can be easily generated. Even larger matrices can be generated by means of interpolation techniques [24, 25].
@@
The solver step has received significantly less attention, since the computational time needed is usually lower. However, if the solver implementation is not optimized, it can become a barrier, preventing the solution of the eigenvalue problem. The issue is particularly severe when the excitonic wave-functions, e.g. the eigenvectors of the BSE, need to be analyzed. The full diagonalization has a scaling $\mathcal{O}(N^{3})$, and would become the most demanding part of the calculation at a certain size. Iterative schemes show a better scaling, typically $\mathcal{O}(N^{2})$. In both approaches, parallel and GPU ported solvers are needed to handle large matrices. Another issue is the storing of the BSE matrix either on disk (taking advantage of parallel I/O) or directly into RAM memory by distributing them over several nodes/GPU cards. In other eigenvalue problems in condensed matter, the matrix can be easily stored in memory and/or even generated on the fly when using iterative solvers, since (i) the generation process is much faster, and (ii) many matrix elements can be related one to the other. The main example in condensed-matter physics is the eigenvalue problem of DFT, where, as opposed to BSE, most efforts have been devoted to develop efficient and highly parallelized iterative algorithms for the solver step, allowing to reach sizes of the order of $10^{7}$ which are not feasible in BSE. In the BSE instead, explicit storage is needed even if iterative solvers are adopted. Also accounting for some memory redundancy used by the available algorithms, this can easily lead to exceeding the barrier of 1 TB of data to be handled. In the most demanding cases, even tens of TB for a single simulation are required.
@@
In the present work, we analyze the performances of two BSE solvers, one based on direct diagonalization, and the other based on an iterative scheme, and we use them to solve prototype BSE matrices. The two approaches have been implemented in the Yambo code, taking advantage of the interfaces with libraries such as ScaLAPACK [26], ELPA [27, 28], PETSc [29] and SLEPc [30]. We consider both the Hermitian BSE, TDA approximation, and the case in which the full structured pseudo-hermitian matrix is used. For the latter case, we discuss in detail how we take advantage of the pseudo-hermitian structure in both solvers to reduce the computational time. We provide performances on parallel cluster, using a local cluster for CPU only simulations, and the Leonardo machine for GPU ported simulations. An analysis of the memory used on the host is also provided.
@@ 2 Description of the algorithm
In this section we provide details about the two computational approaches: full diagonalization and iterative solver. If eigenvectors for more than 5% of the spectrum are required, direct diagonalization routines are recommended. For smaller fractions of the spectrum, iterative solvers are preferable.
2.1 The diagonalization scheme
In this section, we briefly outline the special algorithm for diagonalizing the BSE Hamiltonian given in Eq. (1) when $A=R^{*}$ and $C=C^{T}$, as shown in Ref. [7]. The first step in solving the BSE Hamiltonian $H^{2p}$ with this special structure is to construct a real symmetric matrix $M_{r}$ of the same dimension, which satisfies the relation $Q^{\dagger}H^{2p}Q=-iJM_{r}$, where $J$ is a real skew-symmetric matrix and $Q$ is a unitary matrix:
@@ 2.1 The diagonalization scheme
If $M$ is positive definite, then $M_{r}$ is also positive definite [7], and we can perform a Cholesky decomposition on the real symmetric matrix $M_{r}$, e.g. $M_{r}=LL^{T}$ where $L$ is a real lower triangular matrix. This allows us to cast the original eigenvalue problem into the following form:
@@
where $W=L^{T}JL$ is a real skew-symmetric matrix, and $-iW$ is Hermitian, having the same eigenvalues as the original BSE Hamiltonian. In Eq. (4), the eigenvalues of $W$ differ from the $H^{2p}$ eigenvalues by a factor of $-i$. The eigenvalues of a skew-symmetric matrix come in pairs and are purely imaginary, the eigenvalues of $H^{2p}$ are real and occur in pairs, e.g. $(-\omega_{I},\omega_{I})$. Therefore, the entire BSE problem can be viewed as a real skew-symmetric eigenvalue problem.
@@
In diagonalization, the tridiagonalization step is the computationally expensive part. The algorithm described here enables the tridiagonalization of a real matrix instead of a Hermitian one, significantly reducing both floating-point operations and storage requirements. Once the tridiagonal form of the skew-symmetric matrix is obtained, multiplying the result by $-i$ allows the use of Hermitian tridiagonal solvers to compute the eigenvectors of $W$ [^1] Solving the tridiagonal eigenvalue problem is generally not the primary computational bottleneck, as efficient algorithms are available such as the MRRR method, which scales as $\mathcal{O}(N^{2})$ [31], or the divide-and-conquer approach, which scales as $\mathcal{O}(N^{3})$ but leverages highly optimized General Matrix-Matrix Multiplication (GEMM) operations. After computing the eigenvectors, a back-transformation is applied to retrieve those of the BSE matrix. Moreover, the left eigenvectors for positive eigenvalues, and the left/right eigenvectors corresponding to its negative partner can be determined without explicit computation [7]. More details on the implementation can be found in Ref. [32].
@@
To efficiently interface the algorithm above described with the Yambo code, an MPI redistribution from the Yambo parallel structure to a block-cyclic layout is performed in two main steps. First, the matrix elements are locally grouped along with their corresponding global indices based on their destination rank. Then, a single MPI_AlltoAllv operation is used to transfer the data to their respective destination according to the block-cyclic scheme. Once the matrix is distributed correctly, standard functions from linear algebra libraries can be applied for efficient computations. Here we use ScaLAPACK  [^2] Since ScaLAPACK does not support a real skew-symmetric solver, the tridiagonalization routines for symmetric matrices can be slightly modified to handle skew-symmetric matrices, as both mostly share the same computational operations. and ELPA. For the TDA case, i.e., when the BSE matrix is Hermitian, we use standard Hermitian eigensolvers. ELPA provides two types of Hermitian solver, with type-2 set as the default in Yambo. In ScaLAPACK, the p?heevr solver is used. Both solvers allow one to compute the lowest part of the spectrum. Additionally, ScaLAPACK enables the extraction of eigenvectors within a specified range of eigenvalues or indices.
@@ 2.2 The iterative solver
In many cases, a few eigenvalues are sufficient to calculate the optical absorption spectrum with enough accuracy, so an iterative method can be used. Yambo offers two possible approaches: the first is to use a Lanczos iteration with quadrature rules to obtain the optical absorption spectrum directly, without explicitly computing the eigenpairs. The second approach is to use SLEPc to obtain the eigenpairs, as the eigenvectors provide additional information to compute the wave functions. Prior to the release of SLEPc 3.22, the eigenproblem was solved with a standard non-Hermitian two-sided eigensolver, with the possibility to use shift-and-invert to accelerate the convergence of the desired eigenvalues, which are those closest to zero, and lie in the innermost part of the spectrum because of its symmetry. Since the release of SLEPc 3.22, three new methods are available based on [33, 34], which are three variations of a structure-preserving Lanczos iteration with a thick restart.
@@
Essentially, the problem is projected into a smaller structured problem by building a specific Krylov vector basis, which also has a block structure because of the choice of the initial vector. The basis vectors need to be computed accurately and stored in order to obtain the approximate eigenvectors. To prevent the basis from growing too large, a thick restart is used, taking care not to destroy the structure. These methods exploit the structure and properties of the matrix both numerically (symmetry of the matrix blocks and the spectrum) and computationally, working implicitly with the lower blocks of the matrix and vector basis, without storing them explicitly. Additionally, the upper blocks of the matrix can be shell matrices, a type of matrix that does not store its values explicitly, but instead defines how to perform operations with them; in this way, it is possible to reuse Yambo data structures. All three methods are very similar in terms of operations. We focus our performance analysis on the Shao method, which is the current default solver in SLEPc for BSE problems.
@@
The Shao method does not require the use of the computationally expensive shift-and-invert technique because it computes the eigenvalues of $H^{2}$, folding the spectrum so that the internal eigenvalues become external and converge first. A last step is required to obtain the positive eigenvalues and associated eigenvectors of $H$. The negative eigenvalues and their eigenvectors, and the left eigenvectors, can be obtained trivially from them.

### v2

# Performance in solving the Hermitian and pseudo-Hermitian Bethe-Salpeter equation with the Yambo code

@@ Abstract
We analyze the performance of two strategies in solving the structured eigenvalue problem deriving from the Bethe-Salpeter equation (BSE) in condensed matter physics. The BSE matrix is constructed with the Yambo code, and the two strategies are implemented by interfacing Yambo with the ScaLAPACK and ELPA libraries for direct diagonalization, and with the SLEPc library for the iterative approach. We consider both the Hermitian (Tamm-Dancoff approximation) and pseudo-Hermitian forms, addressing dense matrices of three different sizes. A description of the implementation is also provided, with details for the pseudo-Hermitian case. Timing and memory utilization are analyzed on both CPU and GPU clusters. Our results demonstrate that it is now feasible to handle dense BSE matrices of the order of 105.
Copyright attribution to authors.
This work is a submission to SciPost Physics Codebases.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
where $R$ and $A$ are dense Hermitian matrices, $R=R^{\dagger}$ and $A=A^{\dagger}$, while $C$ is a dense rectangular matrix. Here $R^{\dagger}=(R^{*})^{T}$ denotes conjugate transposition. We further define the $M$ and $\Omega$ matrices as
$\Omega=\begin{pmatrix}I&0\\ 0&-I\end{pmatrix},\quad M=\begin{pmatrix}R&C\\ C^{\dagger}&A\end{pmatrix},$ (2)
which allows us to rewrite Eq. (1) as
$\Omega M\begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}=\omega_{I}\begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}.$ (3)
We refer to the full BSE matrix as $H^{2p}$, and to its structure as pseudo-Hermitian [3, 4, 5, 6]. A matrix is said to be pseudo-Hermitian, if there exists a Hermitian linear operator $\eta$, such that $H^{\dagger}=\eta H\eta^{-1}$ [7]. In the case of BSE, the matrix $H^{2p}$ satisfies the condition with respect to the $\eta=\Omega$, and is thus $\Omega$-pseudo-Hermitian. The BSE Hamiltonian, as shown in Eq. (3), can be expressed as the matrix product of two Hermitian matrices. If the $M$ matrix is positive definite, which can be verified by Cholesky factorization, then the BSE problem reduces to a Hermitian-definite generalized eigenvalue problem with real eigenvalues [3]. Otherwise, it should be treated as a general non-Hermitian eigenvalue problem. Additionally, it can be shown that BSE matrix has a block containing transitions of positive energy (resonant matrix $R$), a block containing transitions of negative energy (anti-resonant matrix $A$), and two blocks which mix positive and negative energy transitions (coupling matrices $C$ and $C^{\dagger}$). Neglecting the coupling term, i.e., setting $C=0$, corresponds to the Tamm-Dancoff approximation (TDA) [1, 8]. Regardless of $M$ being positive definite or not, the BSE matrix in Eq. (1) must have a real eigenspectrum to describe a physical system, provided that lifetimes effects are not included, i.e., all the matrices are frequency independent, as it is usually the case. The structure of the BSE also allows further simplifications of the eigenvalue problem in the special case where R = A, as discussed later. This enables more efficient solution methods that go beyond the standard Hermitian-definite generalized eigenvalue approach [9].
The same structure also appears in the eigenvalue problems derived within Time-Dependent Density Functional Theory (TD-DFT) and within the Random-Phase Approximation (RPA).
@@
The BSE eigenvalue problem is coded in many condensed-matter (and quantum chemistry) packages, both open-source, such as Yambo [10, 11], BerkeleyGW [2, 12], Abinit [13, 14], Exc [15], Exciting [16, 6, 17], GPAW [18, 19], Simple [20], MolGW [21] and commercial or private, such as VASP [22, 5], TurboMole [23], and Fiesta  [24, 25]. Considering also TD-DFT and RPA, the number of scientific codes implementing the eigenvalue problem of Eq. (1) becomes very large, including applications in quantum chemistry and nuclear physics. Most of them use the Hermitian-definite generalized eigenvalue solvers implemented in eigensolver packages. In condensed matter, most BSE applications are done within the TDA, where the BSE eigenvalue problem becomes Hermitian. Indeed, at variance with the case of isolated systems studied in quantum chemistry, the BSE can be recast into a Hermitian quadratic problem only in presence of time-reversal symmetry [5]. However, there are physical cases where going beyond the TDA is needed, and time-reversal is not available [26]. Thus, the full pseudo-Hermitian problem must be addressed [3, 4, 6].
@@
For the case of optical properties, we have $A=R^{*}$, while $C=C^{T}$. The size of both $R$ and $C$ is $N\times N$. Defining $N_{\mathbf{k}}$ as the number of k-points in the full Brillouin zone, $N_{v}$ as the number of included occupied (valence) states, and $N_{c}$ as the number of unoccupied (conduction) states, $N=N_{\mathbf{k}}\,N_{v}\,N_{c}$ is the total number of optical transitions. Numerically, the most demanding step is the building of the $R$ and $C$ matrices. This operation scales as $\mathcal{O}(N^{2})$, with a large prefactor that depends on the square of the basis-set size. For plane-wave codes, this corresponds to $\mathcal{O}(N_{G}^{2})$ where $N_{G}$ is the number of G-vectors. This is why, with the advent of HPC machines and complex structures, a significant effort has been devoted to improve the matrix generation algorithm, taking advantage of MPI, OpenMP and GPU porting schemes such as CUDA Fortran, OpenMP5 or OpenACC. Thanks to these developments, matrices with a size of the order $N\propto 10^{5}$ can be easily generated. Even larger matrices can be generated by means of interpolation techniques [27, 28].
@@
The solver step has received significantly less attention, since the computational time needed is usually lower. However, if the solver implementation is not optimized, it can become a barrier, preventing the solution of the eigenvalue problem. The issue is particularly severe when the excitonic wave-functions, i.e., the eigenvectors of the BSE, need to be analyzed. The full diagonalization has a scaling $\mathcal{O}(N^{3})$, and would become the most demanding part of the calculation at a certain size. Iterative schemes show a better scaling, typically $\mathcal{O}(N^{2})$. In both approaches, parallel and GPU ported solvers are needed to handle large matrices. Another issue is the storing of the BSE matrix, either on disk (taking advantage of parallel I/O) or directly into RAM memory by distributing it over several nodes or GPU cards. In other eigenvalue problems in condensed matter, the matrix can be easily stored in memory and/or even generated on the fly when using iterative solvers, since (i) the generation process is much faster, and (ii) many matrix elements can be related to each other. The main example in condensed-matter physics is the eigenvalue problem of DFT, where, as opposed to BSE, most efforts have been devoted to develop efficient and highly parallelized iterative algorithms for the solver step, allowing to reach sizes of the order of $10^{7}$ which are not feasible in BSE. In the BSE instead, explicit storage is needed even if iterative solvers are adopted. Also accounting for some memory redundancy used by the available algorithms, this can easily lead to exceeding the barrier of 1 TB of data to be handled. In the most demanding cases, even tens of TB for a single simulation are required.
@@
In the present work, we analyze the performances of BSE solvers, two based on direct diagonalization, and the other based on an iterative scheme, and we use them to solve prototype BSE matrices. The two approaches have been implemented in the Yambo code, a plane-wave code, taking advantage of the interfaces with libraries such as ScaLAPACK [29], ELPA [30, 31, 32], PETSc [33] and SLEPc [34]. We consider both the Hermitian BSE within the TDA, and the full structured pseudo-Hermitian BSE matrix. For the latter case, we discuss in detail how we take advantage of the pseudo-Hermitian structure in both solvers to reduce the computational time. We provide parallel performances, using a local cluster for CPU only simulations, and the Leonardo machine for GPU ported simulations. An analysis of the memory used on the host is also provided.
@@ 2 Description of the Algorithm
In this section we provide details about the two computational approaches: full diagonalization and iterative solver.
2.1 The Exact Diagonalization
2.1.1 The Diagonalization Scheme for Full BSE Matrix (Coupling Case)
We briefly outline the special algorithm for diagonalizing the BSE Hamiltonian beyond TDA given in Eq. (1) when $A=R^{*}$ and $C=C^{T}$, as shown in Ref. [9]. The first step in solving the BSE Hamiltonian, $H^{2p}$, with this special structure is to construct a real symmetric matrix $M_{r}$ of the same dimension, which satisfies the relation $Q^{\dagger}H^{2p}Q=-iJM_{r}$, where $J$ is a real skew-symmetric matrix and $Q$ is a unitary matrix given by:
@@ 2.1.1 The Diagonalization Scheme for Full BSE Matrix (Coupling Case)
If $M$ is positive definite, then $M_{r}$ is also positive definite [9], and we can perform a Cholesky decomposition on the real symmetric matrix $M_{r}$, i.e., $M_{r}=LL^{T}$ where $L$ is a real lower triangular matrix. This allows us to cast the original eigenvalue problem into the following form:
@@
where $W=L^{T}JL$ is a real skew-symmetric matrix, and $-iW$ is Hermitian, having the same eigenvalues as the original BSE Hamiltonian. In Eq. (6), the eigenvalues of $W$ differ from the $H^{2p}$ eigenvalues by a factor of $-i$. The eigenvalues of a skew-symmetric matrix come in pairs and are purely imaginary, the eigenvalues of $H^{2p}$ are real and occur in pairs, i.e., $(-\omega_{I},\omega_{I})$. Therefore, the entire BSE problem can be viewed as a real skew-symmetric eigenvalue problem.
@@
In diagonalization, the tridiagonalization step is the computationally expensive part. The algorithm described here enables the tridiagonalization of a real matrix instead of a Hermitian one, significantly reducing both floating-point operations and storage requirements. Once the tridiagonal form of the skew-symmetric matrix is obtained, multiplying the result by $-i$ allows the use of Hermitian tridiagonal solvers to compute the eigenvectors of $W$ [32, 35]. After computing the eigenvectors, a back-transformation is applied to retrieve those of the BSE matrix. Moreover, the left eigenvectors for positive eigenvalues, and the left/right eigenvectors corresponding to its negative partner can be determined without explicit computation [9]. More details on the implementation can be found in Ref. [36]. Consequently, we only need to obtain the right eigenvectors associated with positive eigenvalues. The left and right eigenvectors of the positive and negative eigenvalues for the BSE matrix are related as follows [9]:
$\begin{pmatrix}\text{Right eigenvector for }\omega_{I}&\text{Left eigenvector for }\omega_{I}\\ \begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}&\begin{pmatrix}X_{I}\\ -Y_{I}\end{pmatrix}\\ \text{Right eigenvector for }-\omega_{I}&\text{Left eigenvector for }-\omega_{I}\\ \begin{pmatrix}Y_{I}^{*}\\ X_{I}^{*}\end{pmatrix}&\begin{pmatrix}-Y_{I}^{*}\\ X_{I}^{*}\end{pmatrix}\end{pmatrix}$ (7)
where $\omega_{I}>0$.
Unlike Hermitian-definite generalized eigenvalue solvers, this method reduces floating-point operations for the most computationally demanding steps. Furthermore, when a direct Hermitian-definite generalized eigenvalue solver is used on the BSE matrix, it can destroy the special properties of this matrix and potentially break degeneracies [9], which is highly undesirable in these types of calculations. Aside from improving efficiency, this approach is also robust to such errors.
2.1.2 Interface with ScaLAPACK and ELPA
@@
To efficiently diagonalize the excitonic Hamiltonian in Yambo, both in the Hermitian case and, for the coupling case, using the algorithm described above, an MPI redistribution from the Yambo parallel structure to a block-cyclic layout is performed in two main steps. First, the matrix elements are locally grouped along with their corresponding global indices based on their destination rank. Then, a single MPI_AlltoAllv operation is used to transfer the data to their respective destination according to the block-cyclic scheme. Once the matrix is distributed correctly, standard functions from linear algebra libraries can be applied for efficient computations. Here we use ScaLAPACK [37] and ELPA. For the TDA case, i.e., when the BSE matrix is Hermitian, we use standard Hermitian eigensolvers. ELPA provides two types of Hermitian solver, with type-2 set as the default in Yambo. In ScaLAPACK, the p?heevr solver is used. Both solvers allow one to compute the lowest part of the spectrum. One advantage of ELPA is that its newer versions are specifically optimized for skew-symmetric matrices [32]. This makes ELPA particularly well-suited for solving pseudo-Hermitian or skew-symmetric eigenvalue problems. Implementing such solvers with ELPA is more straightforward, whereas using ScaLAPACK for the same purpose would require additional modifications to handle the skew-symmetric structure. However, ScaLAPACK enables the extraction of eigenvectors within a specified range of eigenvalues or indices.
The interface described above has recently been implemented through the development of a new library [36], which is interfaced via the Yambo code. With this new implementation of the diagonalization solver, capable of computing either the full or partial eigenspectrum, we address the long-standing bottleneck related to the diagonalization of BSE matrices in the Yambo code. Although the code is not yet available in the official Yambo release, it is freely distributed through a recently created fork of the code called Lumen [38, 39].
@@ 2.2 The Iterative Solver
Yambo offers two possible approaches: the first is to use a Lanczos iteration with quadrature rules to obtain the optical absorption spectrum directly, without explicitly computing the eigenpairs. The second approach is to use SLEPc to obtain the eigenpairs, as the eigenvectors provide additional information to compute the wave functions. In many cases, a few eigenvalues are sufficient to calculate the optical absorption spectrum with enough accuracy, so an iterative method can be used.
2.2.1 The SLEPc Solver for the Full BSE Matrix (Coupling Case)
Prior to the release of SLEPc 3.22, the eigenproblem was solved with a standard non-Hermitian two-sided eigensolver, with the possibility to use shift-and-invert to accelerate the convergence of the desired eigenvalues, which are those closest to zero, and lie in the innermost part of the spectrum because of its symmetry. Since the release of SLEPc 3.22, three new methods are available based on [40, 41], which are three variations of a structure-preserving Lanczos iteration with a thick restart.
@@ 2.2.1 The SLEPc Solver for the Full BSE Matrix (Coupling Case)
Essentially, the problem is projected into a smaller structured problem by building a specific Krylov vector basis, which also has a block structure because of the choice of the initial vector. The basis vectors need to be computed accurately and stored in order to obtain the approximate eigenvectors. To prevent the basis from growing too large, a thick restart is used, taking care not to destroy the structure. These methods exploit the structure and properties of the matrix both numerically (symmetry of the matrix blocks and the spectrum) and computationally, working implicitly with the lower blocks of the matrix and vector basis, without storing them explicitly. All three methods are very similar in terms of operations. We focus our performance analysis on the Shao method, which is the current default solver in SLEPc for BSE problems.
@@
The Shao method does not require the use of the computationally expensive shift-and-invert technique because it computes the eigenvalues of $H^{2p}$, folding the spectrum so that the internal eigenvalues become external and converge first. A last step is required to obtain the positive eigenvalues and associated eigenvectors of $H^{2p}$. The negative eigenvalues and their eigenvectors, and the left eigenvectors, can be obtained trivially from them. The details of the SLEPc solvers for the Bethe–Salpeter eigenproblem are available in [42].
2.2.2 MPI Distribution and GPU Porting in SLEPc
In the interface between Yambo and SLEPc two different schemes are available. The first, or shell approach, is based on shell matrices. These are matrices that do not store their values explicitly but instead define how operations on them are carried out. This allows the solver to reuse the existing Yambo data and parallel structures directly. Because Yambo efficiently stores the BSE matrix in memory by exploiting its symmetry properties, this approach requires less memory. However, it comes at the cost of relying on a parallel scheme that is optimized for building of the BSE matrix rather than for solving the eigenvalue problem. Additionally, this approach does not support GPU acceleration. The second approach, called the explicit approach is based on a redistribution of the BSE matrix from the block parallel Yambo structure to the SLEPc structure. In SLEPc, MPI parallelization is based on standard one-dimensional row-wise partitioning of matrices and vectors. This is the scheme used by PETSc, which is the library on which SLEPc is built. In the case of GPU simulations, the underlying local computations are offloaded to the NVIDIA GPU using CUDA (or HIP for AMD GPUs). The GPU porting of the Yambo-SLEPc interface was achieved by defining the redistributed matrix directly on GPU, specifically using the SLEPc matrix type MATDENSECUDA. As in the diagonalization case, this new algorithm is not yet included in the official Yambo release, but it is freely available through a recently developed fork of the code called Lumen [38].